Write the thermochemical equation for the following reactions that occur at constant pressure:

a.Decomposition of mercuric oxide, HgO, to mercury and oxygen gas, requires 87.1 kJ of heat

b.Combustion of hexane, C6H12, to carbon dioxide gas and water releases 3923.9 kJ of heat


1
Expert's answer
2022-04-27T06:18:04-0400

a. 2HgO = 2Hg + O2 - 87.1 kJ ∆H = 87.1 kJ

b. C6H12 + 9O2 = 6CO2 + 6H2O + 3923.9 kJ ∆H = 3923.9 kJ